Calculate the finite area enclosed by y = x ^ {^ {3}} and its tangent line at x = \ frac {-1} {2}.
Answer #1

Ansi but n=1 3 so point P(-1/2 - 1/3 at bant P: here and 3(+)= $; grem y=x, and x = -1 x=-1/2 in y = x3 = (-1/2) = -1/8 at whsolving y = 3x11 and y = x3 = 3x+1) 41°= 34+1 x3 g=**) 4 = 23 4 opber x=-1 cuve (lower curve) 4x 3 X-1 (=/ sol -3x-1= 0, but